finding the depth of a rectangle with p=500ft and width=104ft what is the formula?
Answer by dabanfield(803)

Perimeter (P) of a rectangle is the sum of the two width sides (W) plus the sum of the two depth sides (D).
That is: P = W+W+L+L = (2*W) + (2*D)
Substituting for P and W we have: 500 = (2*104) + 2*D
Simplifying we have: 500 = 208 + 2*D
Subtract 208 from both sides of the equation we have: 292 = 2*D
Dividing both sides of the equaation by 2 we have: 146 = D
In general:
P = (2+W) + (2*D)
P-(2*W) = 2*D
(P-(2*W))/2 = D